This is a python questionThe distance between 2 points (xi.Vi). (x2.y2) on a plane is given by the following equation: Write a function named shortest
Answer #1

CODE:

import math
anylist=[[1,1],[2,2],[4,4]] #input list
def shortestDist(anylist):
i=0
j=i+1
z=0
templist=[] #temporary list to store distances
while(z<len(anylist)-1):
sum1=(anylist[i][0]-anylist[j][0] )**2 + ( anylist[i][1]-anylist[j][1] )**2
i=i+1
j=i+1
z=z+1
templist.append(math.sqrt(sum1))
return templist
calling=shortestDist(anylist)
print(min(calling)) # printing minimum distance